solve for r: v= 1/3*pi*r^2h
Answer by harpazo(655)


solve for r:
v = 1/3*pi*r^2h
v = (pir^2h)/3
3v = (pir^2h)
3v/pih = r^2
Take the square of both sides.
SQRT{3v/pih} = SQRT{r^2}
SQRT{3v/pih} = r
Note: SQRT = square root